According to a survey, 38% of employees working at a very large company are happy with their jobs. Suppose that two employees are selected at random from this company. Let x denote the number of employees in this sample of two who are happy with their jobs. Construct the probability distribution table of x.
_______
x p(x)
_______
0 ?
1 ?
2 ?
_______

Answer by ikleyn(52852)

.
According to a survey, 38% of employees working at a very large company are happy with their jobs.
Suppose that two employees are selected at random from this company.
Let x denote the number of employees in this sample of two who are happy with their jobs.
Construct the probability distribution table of x.
_______
x p(x)
_______
0 ?
1 ?
2 ?
_______
~~~~~~~~~~~~~~~~~~~~~~

(a)  x = 0  means that in the pair no one is happy.
     
     The probability for it is  P(x=0) = (1-0.38)*(1-0.38) = 0.3844.



(b)  x = 1  means that one of the two in the pair is happy, while another is not.

     The probability for it is  P(x=1) = (1-0.38)*0.38 + 0.38*(1-0.38) = 0.4712.



(c)  x = 2  means that both in the pair are happy.

     The probability for it is  P(x=2) = 0.38*0.38 = 0.1444.

Solved.